suppose f(z) =1/z. write f in the form f(z) = u(x,y) + iv(x,y), where z = x+iy and u and v are real-valued functions


1
Expert's answer
2022-05-06T05:40:33-0400

"z=x+iy", where "x,y\in{\mathbb{R}}". We receive: "\frac{1}{z}=\frac{1}{x+iy}=\frac{x-iy}{(x+iy)(x-iy)}=\frac{x-iy}{x^2+y^2}." From the latter expression we receive that "u=\frac{x}{x^2+y^2}" and "v=\frac{-y}{x^2+y^2}".

Answer: "f(z)=u(x,y)+iv(x,y)," where "z=x+iy", "x,y\in{\mathbb{R}}"; "u(x,y)=\frac{x}{x^2+y^2}", "v(x,y)=\frac{-y}{x^2+y^2}".